Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

a rare on GMAT

Expert replies
by sanju09 » Tue Mar 02, 2010 4:22 am
10 different books and 2 different pens are given to 3 boys so that each boy gets equal number of things. What is the probability that same boy doesn't receive both pens?
(A) 3/11
(B) ¼
(C) 4/11
(D) 8/11
(E) 9/11
The mind is everything. What you think you become. –Lord Buddha



Sanjeev K Saxena
Quantitative Instructor
The Princeton Review - Manya Abroad
Lucknow-226001

www.manyagroup.com
Join the discussion
Source: — Problem Solving |

by rohan_vus » Tue Mar 02, 2010 4:38 am
Probability that one boy receives both pens is 10c2*2c2/12c4 = 1/11
So total probability that among 3 boys, anyone will have boths pens, is 3*1/11 = 3/11

So probability that none will have both pens = 1 - 3/11 = 8/11
Join the discussion

by stufigol » Wed Mar 03, 2010 12:13 am
u explain the 10c2 and 12c4 plz
thanks
Join the discussion

by harsh.champ » Wed Mar 03, 2010 4:27 am
stufigol wrote:u explain the 10c2 and 12c4 plz
thanks
Ok.
Lets look at rohan's explanation.
Probability that one boy receives both pens is 10c2*2c2/12c4 = 1/11
So total probability that among 3 boys, anyone will have boths pens, is 3*1/11 = 3/11

So probability that none will have both pens = 1 - 3/11 = 8/11
Suppose the same boy gets both the pens.
Now,10C2 = No. of ways in which the 2 books can be selected.
12C4 = Selection of any 4 items out of the (10 + 2) items

Also,the boy can be selected in 3C1 ways.
So, we have our case (10C2 x 2C2 x 3C1)/12c4 = 3/11
Now, P(favorable case) = 1 - P(all the unfavorable cases)

I hope it is clear now. :)
It takes time and effort to explain, so if my comment helped you please press Thanks button :)



Just because something is hard doesn't mean you shouldn't try,it means you should just try harder.

"Keep Walking" - Johnny Walker :P
Join the discussion